OpenStudy (anonymous):

CORRECT ANSWER GETS MEDAL! Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

(1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (anonymous):

some one please?

OpenStudy (debbieg):

Try each point in both inequalities. The one for which the inequality is true, is the winner.

OpenStudy (debbieg):

I'll get you started: for the first point, you have: (1, 10) OK, wait... sure you meant to say "less than or equal to" for each of your inequality statements?? "less than or greater to" makes no sense, so I'm going to assume that isn't what you meant. So you have: \(y \le x - 5\\ y \le -x - 4 \) Trying (1,10) in the first one: if x=1, y=10, you get: \(10 \le 1 - 5 \) so \(10 \le -4 \) which is clearly FALSE, so no need to check the point in the other inequality, that point is not in the solution set. Now try that with the other points, until you find one that makes BOTH inequalities TRUE.